OpenStudy (anonymous):

A train starts from rest moves with a constant acceleration of 2m/s^2 for half a minute. The brakes are then applied and then train comes to rest in one minute after applying breaks. (c) the position(s) of the train at half the maximum speed

OpenStudy (abhisar):

O_o I thought i was tagged !

OpenStudy (anonymous):

Yes u were

OpenStudy (abhisar):

Same process !

OpenStudy (anonymous):

I am not getting the correct answer

OpenStudy (anonymous):

Max speed = 60

OpenStudy (anonymous):

that is correct

OpenStudy (abhisar):

yes

OpenStudy (anonymous):

Now how to proceed

OpenStudy (abhisar):

Use (30)\(^2\) = 0 + 2*2*s

OpenStudy (abhisar):

Find s

OpenStudy (anonymous):

YEs man got it!

OpenStudy (anonymous):

why can't we use s= ut +1/2 at^2

OpenStudy (abhisar):

u can but first u will have to calculate t

OpenStudy (abhisar):

by using V=U+at

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

thanks!!!!

OpenStudy (abhisar):

so it's better to use V^2=U^2+2as

OpenStudy (anonymous):

Don't you think openstudy , wastes much time

OpenStudy (abhisar):

how ?

OpenStudy (anonymous):

I am doing this problem for last 35 minutes or so , and it was rather a simple one , and i have thousands yet to do

OpenStudy (anonymous):

This medium is not useful for bigger problems

OpenStudy (anonymous):

unless u have a lot of time

OpenStudy (anonymous):

what will be the accelration at the instant when it is at maximum speed

OpenStudy (abhisar):

That's true somehow....there are some input limitations...! but if u have a touchscreen then u can do it faster by using drawing tool

OpenStudy (anonymous):

Still , one problem i posted other day took 2.5 hours

OpenStudy (anonymous):

and i didn't realise

OpenStudy (anonymous):

Time flies away when u r on OS

OpenStudy (abhisar):

I'll suggest its better to solve by urself..post them on os at extreme cases. :D

OpenStudy (abhisar):

btw inst accn at the time of max speed will be 2m/s^2

OpenStudy (anonymous):

is it , i framed this question

OpenStudy (abhisar):

Since the train moves with const accn

OpenStudy (abhisar):

yeah it will 2m/s^2

OpenStudy (anonymous):

let's check by equation

OpenStudy (anonymous):

v=u+at 60 =(30)a yes 2 hehe

OpenStudy (anonymous):

sorry for trouble

OpenStudy (abhisar):

LOL ! It's ok :)

OpenStudy (abhisar):

btw u saw the webpage i developed ?

OpenStudy (anonymous):

nopess